1. Consider the basis S = {v1, v2} for R2, where v1 = (− 2, 1) and v2 = (1, 3), and let T:R2 → R3 be the linear transformation such that

T(v1) = (− 1, 2, 0) and T(v2) = (0, − 3, 5)

Find a formula for T(x1, x2), and use that formula to find T(2, − 3).

Give exact answers in the form of a fraction.

2. Consider the basis S = {v1, v2, v3} for R3, where v1 = (1, 1, 1), v2 = (1, 1, 0), and v3 = (1, 0, 0), and let T:R3 → R3 be the linear operator for which

T(v1) = (3, − 1, 6), T(v2) = (4, 0, 1), T(v3) = (− 1, 7, 1)

Find a formula for T(x1, x2, x3), and use that formula to find T(3, 6, − 1).

3. Let v1, v2, and v3 be vectors in a vector space V, and let T:V → R3 be a linear transformation for which

T(v1) = (1, − 1, 2), T(v2) = (0, 3, 2), T(v3) = (− 3, 1, 2)

Find T(4v1 − 5v2 + 6v3).
1
Expert's answer
2020-01-26T16:40:24-0500

"T" - linear transformation:

"T(u+v)=T(u)+T(v)"

"T(ku)=kT(u)"


1.

"v=(x_1,x_2) \in \mathbb{R} ^2" - arbitrary vector.

It can be expressed as a linear combination of "v_1" and "v_2":

"v=av_1+bv_2"

"(x_1,x_2)=a(-2,1)+b(1,3)=(-2a+b,a+3b)"

"\begin{cases}
x_1=-2a+b \\ x_2=a+3b
\end{cases}
\; \;
\begin{cases}
x_1+2x_2=7b\\ x_2=a+3b
\end{cases}
\begin{cases}
b=\frac{x_1}{7}+\frac{2x_2}{7} \\a=\frac{-3x_1}{7}+\frac{x_2}{7}
\end{cases}"


"T(v)=T(av_1+bv_2)=aT(v_1)+bT(v_2)=a(-1,2,0)+b(0,-3,5)="

"=(\frac{-3x_1}{7}+\frac{x_2}{7})(-1,2,0)+(\frac{x_1}{7}+\frac{2x_2}{7} )(0,-3,5)="

"=(\frac{3x_1}{7}+\frac{-x_2}{7} , \frac{-9x_1}{7}+\frac{-4x_2}{7} , \frac{5x_1}{7}+\frac{10x_2}{7})=
\frac{1}{7}(3x_1-x_2,-9x_1-4x_2,5x_1+10x_2)"


"T(2,-3)=\frac{1}{7}(9,-6,-20)=(\frac{9}7,-\frac{6}7,-\frac{20}7)"


2.

"v=(x_1,x_2,x_3)\in \mathbb{R}^3" - arbitrary vector

It can bee expressed as a linear combination of "v_1,v_2,v_3:"

"v=av_1+bv_2+cv_3"

"(x_1,x_2,x_3)=a(1,1,1)+b(1,1,0)+c(1,0,0)=(a+b+c,a+b,a)"

"\begin{cases}
x_1=a+b+c \\
x_2=a+b \\
x_3=a
\end{cases}

\begin{cases}
a=x_3 \\ b=x_2-x_3 \\ c= x_1-x_2
\end{cases}"


"T(v)=T(av_1+bv_2+cv_3)=aT(v_1) +bT(v_2)+cT(v_3)="

"=a(3,-1,6)+b(4,0,1)+c(-1,7,1)="

"= x_3(3,-1,6)+(x_2-x_3)(4,0,1)+(x_1-x_2)(-1,7,1)="

"=(-x_1+5x_2-x_3,7x_1-7x_2-x_3, x_1+5x_3)"

"T(3,6,-1)=(28,-20,-2)"


3.

"T(4v_1-5v_2+6v_3)=4T(v_1)-5T(v_2)+6T(v_3)="

"=4(1,-1,2)-5(0,3,2)+6(-3,1,2)=(-14,-13,10)"